OpenStudy (anonymous):

In a simple random sample of 219 students at a college, 73 reported that they have at least $1000 of credit card debt. What is the 99% confidence interval for the percent of all the students at that college who have at least $1000 in credit card debt? (Points : 3) (25.0, 41.6) (27.5, 39.1) (30.1, 36.5) (31.0, 35.6)

OpenStudy (anonymous):

i need help with how to solve this step by step

OpenStudy (anonymous):

thats a waste of own bucks

OpenStudy (anonymous):

owl

OpenStudy (anonymous):

any ideas?

OpenStudy (michele_laino):

I'm sorry I don't know your answer. Please wait, I ask to another tutor

OpenStudy (michele_laino):

@CGGURUMANJUNATH can you help here please?

OpenStudy (anonymous):

thank you

OpenStudy (dan815):

Hi there :)

OpenStudy (dan815):

Are you familiar with how to use Z tables

OpenStudy (dan815):

you need to first identity, what Z score gives you a 99% confidence interval

OpenStudy (anonymous):

2.567?

OpenStudy (dan815):

sounds about right, ill have to check myself too but, now whats left for you is to find the true standard deviation

OpenStudy (asad786):

dan can you help me on a question after this plz

OpenStudy (anonymous):

I'm getting stuck at 73 +- 28.8

OpenStudy (dan815):

the ratio 73/219 is the sample percentage that have atleast 1000 dollars of debt 73/219 = 33.33%

OpenStudy (dan815):

that mean is the same for the actual distribution as well

OpenStudy (dan815):

you will be using the formula 33.33 +- (2.57*sigma) you need to find the sigma which is the true standard deiviation

OpenStudy (dan815):

also if you notice the choices given to you the middle values are all 33.33 in that range

OpenStudy (dan815):

(25.0 + 41.6)/2 =33.33 (27.5 + 39.1)/2=33.33 (30.1 + 36.5)/2 = 33.33 (31.0, + 35.6)/2 = 33.33

OpenStudy (badmood):

~clapping~

OpenStudy (anonymous):

B?

OpenStudy (dan815):

uhh wait im thinking how to get the true STD from a mean and sample size

OpenStudy (anonymous):

:s

OpenStudy (anonymous):

so would the answer be B?

OpenStudy (dan815):

Show me some of your work, maybe ill see if it makes sense at least

OpenStudy (anonymous):

the answer is A

OpenStudy (anonymous):

Ok in order to find the sample standard deviation we need to realize that we are using proportions here. Since the answer gives the percentage as a whole number for computational sake we will do the same. p=73/219 =33% q=(1-p)=1-33=67% We know that our sample standard dev is \( \sqrt{ \frac{p*q}{n}}\) So S=\( \sqrt{ \frac{33*67}{216}}=3.17\) Therefore our confidence interval is \( 33.33 \pm 2.58*3.17 \to (25.15,41.5)\)